a businessman requires 2800 kcal energy daily. how many glucose and atp does he require to produce this much energy?
 

Grade:10

1 Answers

Prajwal Chauhan
8 years ago
1 ATP = 7.3 kcal
so, no. of ATP molecules = 2800/7.3 = 383.56
and 1 molecule of glucose provides 38 molecules of ATP ( glycolysis,kreb’s cycle, etc.)
so glucose required will be = 383.56/38 = 10.09